OpenStudy (questionaskerduh1243):

Will fan and medal \ Write the number of permutations in factorial form. then simplify how many diffrent ways can you and five friends sit in your assigned seats when you go to a concert

satellite73 (satellite73):

in a row?

OpenStudy (questionaskerduh1243):

uh im not sure you want to see the multiple answers to help?

satellite73 (satellite73):

\[\huge 5!=5\times 4\times 3\times 2\]